# Thermodynamics of charmed hadrons across chiral crossover from lattice QCD

### v1

@@ Abstract
We use up to fourth-order charm fluctuations and their correlations with net baryon number, electric charge, and strangeness fluctuations, calculated within the framework of lattice QCD, to study the continuum partial pressure contributions of charmed baryons and mesons. We show that, at and below the chiral crossover temperature, these partial pressures receive enhanced contributions from experimentally unobserved charmed hadrons predicted by the Quark Model. Additionally, we demonstrate that at the chiral crossover, the Hadron Resonance Gas description breaks down, signaling the onset of charm deconfinement. We present evidence for the survival of low-lying non-radial $1S$ and $1P$ hadron-like excitations above the chiral crossover. Finally, we investigate the continuum partial pressure contribution of charm quark-like excitation that emerges at the chiral crossover and calculate its temperature-dependent in-medium mass.
@@ 1 Introduction
The chiral crossover also appears as the charm deconfinement transition, i.e. as the onset of charm quarks as new degrees of freedom Bazavov et al. (2024) . Partial charm hadronic pressures start to deviate from the HRG model prediction, but not all charmed hadrons dissociate at $T_{pc}$. Furthermore, recent lattice QCD studies provide compelling evidence for the survival of charmed hadrons above the chiral crossover temperature Bazavov et al. (2024); Sharma (2024a, b). In this paper, we investigate the gradual deconfinement of charmed hadrons by analyzing continuum extrapolated proxies for partial pressures of charmed baryons and mesons across the chiral crossover.
@@ 2 Calculating generalized susceptibilities on the lattice
where $\vec{\mu}=(\mu_{B},\mu_{Q},\mu_{S},\mu_{C})$. Note that ${\chi^{BQSC}_{klmn}}$ will be non-zero only for ${(k+l+m+n)\in\text{even}}$. In the following, if the subscript corresponding to a conserved charge is zero in the left hand side of Eq. 2, then both the corresponding superscript as well the zero subscript will be suppressed.
@@ 3 Boltzmann Approximation
Figure 1: Comparision of QM-HRG and PDG-HRG predictions for charmed baryons [Left] and charmed mesons [Right]. $\Delta=100(1-\text{QM-HRG}/\text{PDG-HRG})|_{T_{pc}}$. The yellow bands represent $T_{pc}$ with its uncertainty.
@@ 3.1 Charmed Hadrons
Hadron resonance gas model (HRG) has been successful in describing the particle abundance ratios measured in the heavy-ion experiments. It describes a non-interacting gas of hadron resonances, and therefore can be used to calculate the hadronic pressure below ${T_{pc}}$ Andronic et al. (2019). In the Boltzmann approximation, the dimensionless partial pressures from the charmed meson, ${M_{C}},$ and the charmed baryon, ${B_{C}}$, sectors take the following forms Allton et al. (2005):
In above equations, at a given temperature, the summation is over all charmed mesons/baryons (C-mesons/baryons) with masses given by ${m_{i}}$; degeneracy factors of the states with equal mass and same quantum numbers are represented by ${g_{i}}$; ${K_{2}(x)}$ is a modified Bessel function, which for a large argument can be approximated by ${K_{2}(x)}\sim\sqrt{\pi/2x}\;e^{-x}\;[1+\mathbb{O}(x^{-1})]$. Consequently, if a charmed state under consideration is much heavier than the relevant temperature scale, such that ${m_{i}\gg T}$, then the contribution to ${P_{C}}$ from that particular state will be exponentially suppressed, e.g., the singly-charmed ${\Lambda}_{c}^{+}$ baryon has a PDG mass of about $2286$ MeV, whereas the doubly-charmed ${\Xi_{cc}^{++}}$ baryon’s mass as tabulated in PDG records is about $3621$ MeV, hence at ${T_{pc}}$, the contribution to ${B_{C}}$ from ${\Xi_{cc}^{++}}$ will be suppressed by a factor of $10^{-4}$ in relation to ${\Lambda}_{c}^{+}$ contribution. Therefore, in the validity regime of Boltzmann approximation, substituting Eq. LABEL:eq:McBc in Eq. 2 implies $P_{C}(T,\vec{\mu})\approx\chi_{2}^{C}\approx\chi_{n}^{C}$, for $n$ even.
Charm fluctuations calculated in the framework of lattice QCD can receive enhanced contributions due the existence of not-yet-discovered open-charm states. It is possible to compare this enhancement to the HRG calculations performed with two data sets. The first scenario, denoted by PDG-HRG, is based on the experimentally established states tabulated in the PDG records. The second scenario, denoted by QM-HRG, in addition to PDG states, takes into account states predicted via Quark-Model calculations Ebert et al. (2010, 2011); Chen et al. (2023). In the validity regime of HRG, $\chi^{nm}_{BC}\approx B_{C}$, and $M_{C}=P_{C}-B_{C}$. Fig. 1 [Left] shows that the QM-HRG predicts a $92\%$ enhancement of the partial charmed baryon pressure at $T_{pc}$ compared to the PDG-HRG expectation, whereas the predicted enhancement in the partial charmed meson pressure is only $15\%$. These results suggest that, based on the QM-HRG predictions, the charmed baryonic sector is significantly more incomplete than the charmed mesonic sector. In recent years, the experimental discovery of various $\Lambda_{C}$ and $\Omega_{C}$ resonances, in particular by LHCb and Belle collaborations Kato et al. (2016); Aaij et al. (2017a, b); Chen and Liu (2017) have confirmed this.
@@ 3.2 Charm Quarks
Charm quarks offer an advantage over the light quarks because for temperatures a few times $T_{pc}$, the Boltzmann approximation works for an ideal massive quark-antiquark gas Allton et al. (2005); Bazavov et al. (2014). Therefore, in this approximation, the dimensionless partial charm quark pressure, $Q_{C}$ is given by,
$Q_{C}(T,\overrightarrow{\mu})=\dfrac{3}{\pi^{2}}\bigg{(}\dfrac{m_{q}^{C}}{T}\bigg{)}^{2}K_{2}(m_{q}^{C}/T)\text{cosh}\bigg{(}\dfrac{2}{3}\hat{\mu}_{Q}+\dfrac{1}{3}\hat{\mu}_{B}+\hat{\mu}_{C}\bigg{)}\text{ ,}$ (3)
@@ 4 Enhancement of charmed hadrons
Continuum estimates for charmed baryon and meson partial pressures shown in Fig. 2 are obtained by multiplying continuum extrapolated $\chi^{4}_{C}$ values to ratios $P^{C}_{B}/P_{C}$ and $P^{C}_{M}/P_{C}$ shown in our previous work Bazavov et al. (2024). As mentioned earlier, these ratios, despite being calculated on a finite temporal lattice extent, are largely cut-off independent. For $T<T_{pc}$, Fig. 2 shows a clear agreement of the QM-HRG predictions with the partial hadron pressures calculated on the lattice. This agreement, in addition to corroborating the validity of HRG in the low temperature phase, confirms the existence of experimentally unobserved hadronic states. In particular, at the chiral crossover, in the charmed baryonic sector, $P^{C}_{B}$ calculated on the lattice is almost twice as large as the PDG-HRG expectation, whereas $P^{C}_{M}$ calculated on the lattice is around $20\%$ larger than the PDG-HRG prediction. The enhancement factors explicitly quoted in Fig. 2 agree within errors with the enhancement predictions quoted in Fig. 1.
@@ 5 Charm degrees of freedom in QGP
Departure of charmed hadronic pressures from their respective QM-HRG predictions hints at a sequential melting pattern. This means that the higher excited charmed hadrons start dissociating at $T_{pc}$, whereas the ground state charmed hadrons survive inside the QGP. To put this on more solid grounds, we performed another HRG model calculation based on the low-lying, $1S$ and $1P$, charmed hadrons tabulated in Chen et al. (2023) – we label this as 1S1P-HRG. For $T_{pc}<T\leq 166.1$ MeV, both $P^{C}_{B}$ and $P^{C}_{M}$ in Fig. 2 are described by 1S1P-HRG. For $P^{C}_{M}$, both PDG-HRG and 1S1P-HRG almost overlap, and only at the highest temperatures shown in Fig. 2 [Right] there are some visible differences. This is expected because all low-lying charmed meson states are experimentally known. In addition to low-lying 1S and 1P states, PDG-HRG contains one charmed-light meson: $D_{3}^{*}(2750)$ and two charmed-strange mesons: $D_{s1}^{*}(2700)$ and $D_{s3}^{*}(2860)$. We used PDG masses of $D$, $D^{*}$, $D^{*}_{0}$ and $D_{1}$ charmed-light mesons. If chiral partners were to get degenerate at $T_{pc}$, then the scalar and axial-vector mesons would get lighter, making 1S1P-curve curve overshoot $P^{C}_{M}$ above $T_{pc}$. This indicates that the charm sector is not influenced by the chiral symmetry restoration. Similar observation holds for the charmed-strange mesonic sector. In the charmed baryonic sector, after departure from 1S1P-HRG, for the highest two temperatures, PDG-HRG prediction can very well describe $P^{C}_{B}$ in Fig. 2 [Left]. The fact that 1S1P-HRG lies above PDG-HRG for $P^{C}_{B}$ indicates that there are many low-lying and therefore thermodynamically dominant charmed baryons missing from the PDG record.
In the coexistence phase of charmed hadron and quark-like excitations, $m^{C}_{q}$ in Eq. 3 becomes temperature dependent and can be interpreted as the mass of a quasi-particle with quantum numbers of charm quark. Given $P^{C}_{q}(T)$ as a function of temperature, one can solve Eq. 3 at each temperature and obtain temperature dependence of $m^{C}_{q}$. In Fig. 3 [Right], at each temperature, the error on $m^{C}_{q}$ is the standard deviation of $m^{C}_{q}$ values obtained after solving Eq. 3 for $50$ fake Gaussian samples. Fig. 3 [Right] shows that at $T=162.4$ MeV, $m^{C}_{q}$ is around the mass of D-meson, and starts decreasing with temperature. Quasi-particle model in Eq. 4 considers a non-interacting gas of charmed hadron and quark-like excitations, but the temperature dependence of $m^{C}_{q}$ encodes the in-medium interactions. In the non-interacting quark gas limit, $m^{C}_{q}$ will become the pole mass of charm quark.

### v2

@@ Abstract
We use up to fourth-order charm fluctuations and their correlations with net baryon number, electric charge, and strangeness fluctuations, calculated within the framework of lattice QCD, to study the continuum partial pressure contributions of charmed baryons and mesons. We show that, at and below the chiral crossover temperature, these partial pressures receive enhanced contributions from experimentally unobserved charmed hadrons predicted by the Quark Model. Additionally, we demonstrate that at the chiral crossover, the Hadron Resonance Gas description breaks down, signaling the modification of open charm hadrons and thereby implying the onset of charm deconfinement. We present evidence for the survival of low-lying non-radial $1S$ and $1P$ hadron-like excitations above the chiral crossover, which hints at the sequential melting of charmed hadrons. Finally, we investigate the continuum partial pressure contribution of charm quark-like excitation that emerges at the chiral crossover and calculate its temperature-dependent in-medium mass.
@@ 1 Introduction
The chiral crossover also exhibits deconfining feature in the charm quark sector, i.e. charm quarks as new degrees of freedom start appearing at or close to the pseudo-critical temperature of chiral symmetry restoration (Bazavov et al. (2024)). Clearly the appearance of charm quark degrees of freedom at $T_{pc}$, arising due to the disappearance of some hadronic states, will also liberate light quark degrees of freedom, although in the chiral limit this may not directly be related to the universal aspects of the chiral phase transition. The partial charm hadronic pressures start to deviate from the HRG model prediction, but not all charmed hadrons dissociate at $T_{pc}$. Furthermore, recent lattice QCD studies provide compelling evidence for the survival of charmed hadrons above the chiral crossover temperature Bazavov et al. (2024); Sharma (2024a, b). In this paper, we investigate the gradual deconfinement of charmed hadrons by analyzing continuum extrapolated proxies for partial pressures of charmed baryons and mesons across the chiral crossover.
@@ 2 Calculating generalized susceptibilities on the lattice
where $\vec{\mu}=(\mu_{B},\mu_{Q},\mu_{S},\mu_{C})$. Note that ${\chi^{BQSC}_{klmn}}$ will be non-zero only for ${(k+l+m+n)\in\text{even}}$. In the following, if the subscript corresponding to a conserved charge is zero in the left hand side of Eq. (2), then both the corresponding superscript as well the zero subscript will be suppressed.
@@ 3 Boltzmann Approximation
Figure 1: Comparison of QM-HRG and PDG-HRG predictions for charmed baryons [Left] and charmed mesons [Right]. $\Delta=100|1-\text{QM-HRG}/\text{PDG-HRG}|_{T_{pc}}$. The yellow bands represent $T_{pc}$ with its uncertainty.
@@ 3.1 Charmed Hadrons
Hadron resonance gas model (HRG) has been successful in describing the particle abundance ratios measured in the heavy-ion experiments. It describes a non-interacting gas of hadron resonances, and therefore can be used to calculate the hadronic pressure below ${T_{pc}}$ Andronic et al. (2019). In the Boltzmann approximation, the dimensionless partial pressures from the charmed meson, ${P^{C}_{M}},$ and the charmed baryon, ${P^{C}_{B}}$, sectors take the following forms Allton et al. (2005):
In above equations, at a given temperature, the summation is over all charmed mesons/baryons (C-mesons/baryons) with masses given by ${m_{i}}$; degeneracy factors of the states with equal mass and same quantum numbers are represented by ${g_{i}}$; ${K_{2}(x)}$ is a modified Bessel function, which for a large argument can be approximated by ${K_{2}(x)}\sim\sqrt{\pi/2x}\;e^{-x}\;[1+\mathbb{O}(x^{-1})]$. Consequently, if a charmed state under consideration is much heavier than the relevant temperature scale, such that ${m_{i}\gg T}$, then the contribution to ${P_{C}}$ from that particular state will be exponentially suppressed, e.g., the singly-charmed ${\Lambda}_{c}^{+}$ baryon has a PDG mass of about $2286$ MeV, whereas the doubly-charmed ${\Xi_{cc}^{++}}$ baryon’s mass as tabulated in PDG records is about $3621$ MeV, hence at ${T_{pc}}$, the contribution to ${P^{C}_{B}}$ from ${\Xi_{cc}^{++}}$ will be suppressed by a factor of $10^{-4}$ in relation to ${\Lambda}_{c}^{+}$ contribution. Therefore, in the validity regime of Boltzmann approximation, substituting Eq. (LABEL:eq:McBc) in Eq. (2) implies $P_{C}(T,\vec{\mu})\approx\chi_{2}^{C}\approx\chi_{n}^{C}$, for $n$ even.
Charm fluctuations calculated in the framework of lattice QCD can receive enhanced contributions due the existence of not-yet-discovered open-charm states. It is possible to compare this enhancement to the HRG calculations performed with two data sets. The first scenario, denoted by PDG-HRG, is based on the experimentally established states tabulated in the PDG records. The second scenario, denoted by QM-HRG, in addition to PDG states, takes into account states predicted via Quark-Model calculations Ebert et al. (2010, 2011); Chen et al. (2023). In the validity regime of HRG, $\chi^{nm}_{BC}\approx P^{C}_{B}$, and $P^{C}_{M}=P_{C}-P^{C}_{B}$. Fig. 1 [Left] shows that the QM-HRG predicts a $92\%$ enhancement of the partial charmed baryon pressure at $T_{pc}$ compared to the PDG-HRG expectation, whereas the predicted enhancement in the partial charmed meson pressure is only $15\%$. These results suggest that, based on the QM-HRG predictions, the charmed baryonic sector is significantly more incomplete than the charmed mesonic sector. In recent years, the experimental discovery of various $\Lambda_{C}$ and $\Omega_{C}$ resonances, in particular by LHCb and Belle collaborations Kato et al. (2016); Aaij et al. (2017a, b); Chen and Liu (2017) have confirmed this.
@@ 3.2 Charm Quarks
Charm quarks offer an advantage over the light quarks because for temperatures a few times $T_{pc}$, the Boltzmann approximation works for an ideal massive quark-antiquark gas Allton et al. (2005); Bazavov et al. (2014). Therefore, in this approximation, the dimensionless partial charm quark pressure, $P^{C}_{q}$, is given by,
$P^{C}_{q}(T,\overrightarrow{\mu})$ $=$ $\dfrac{3}{\pi^{2}}\bigg{(}\dfrac{m_{q}^{C}}{T}\bigg{)}^{2}K_{2}(m_{q}^{C}/T)\cdot$ (3)
$\text{cosh}\bigg{(}\dfrac{2}{3}\hat{\mu}_{Q}+\dfrac{1}{3}\hat{\mu}_{B}+\hat{\mu}_{C}\bigg{)}\text{ ,}$
@@ 4 Enhancement of charmed hadrons
Continuum estimates for charmed baryon and meson partial pressures shown in Fig. 2 are obtained by multiplying continuum extrapolated $\chi^{C}_{4}$ values to ratios $P^{C}_{B}/P_{C}$ and $P^{C}_{M}/P_{C}$ shown in our previous work Bazavov et al. (2024). As mentioned earlier, these ratios, despite being calculated on a finite temporal lattice extent, are largely cut-off independent. For $T<T_{pc}$, Fig. 2 shows a clear agreement of the QM-HRG predictions with the partial hadron pressures calculated on the lattice. This agreement, in addition to corroborating the validity of HRG in the low temperature phase, confirms the existence of experimentally unobserved hadronic states. In particular, at the chiral crossover, in the charmed baryonic sector, $P^{C}_{B}$ calculated on the lattice is almost twice as large as the PDG-HRG expectation, whereas $P^{C}_{M}$ calculated on the lattice is around $20\%$ larger than the PDG-HRG prediction. The enhancement factors explicitly quoted in Fig. 2 agree within errors with the enhancement predictions quoted in Fig. 1.
@@ 5 Charm degrees of freedom in QGP
Departure of charmed hadronic pressures from their respective QM-HRG predictions hints at a sequential melting pattern. This means that the higher excited charmed hadrons start dissociating at $T_{pc}$, whereas the ground state charmed hadrons survive inside the QGP. To put this on more solid grounds, we performed another HRG model calculation based on the low-lying, $1S$ and $1P$, charmed hadrons tabulated in Chen et al. (2023) – we label this as 1S1P-HRG. For $T_{pc}<T\leq 166.1$ MeV, both $P^{C}_{B}$ and $P^{C}_{M}$ in Fig. 2 are described by 1S1P-HRG. For $P^{C}_{M}$, both PDG-HRG and 1S1P-HRG almost overlap, and only at the highest temperatures shown in Fig. 2 [Right] there are some visible differences. This is expected because all low-lying charmed meson states are experimentally known. In addition to low-lying 1S and 1P states, PDG-HRG contains one charmed-light meson: $D_{3}^{*}(2750)$ and two charmed-strange mesons: $D_{s1}^{*}(2700)$ and $D_{s3}^{*}(2860)$. We used PDG masses of $D$, $D^{*}$, $D^{*}_{0}$ and $D_{1}$ charmed-light mesons. If chiral partners were to get degenerate at $T_{pc}$, then the scalar and axial-vector mesons would get lighter, making 1S1P-HRG curve overshoot $P^{C}_{M}$ above $T_{pc}$. This indicates that the charm sector is not strongly influenced by the chiral symmetry restoration. This is also in line with studies of screening masses for charmed parity partners, which indicate that unlike in the light quark sector the charmed screening masses become almost degenerate only at temperatures above $2T_{pc}$ Bazavov et al. (2015). Similar observation holds for the charmed-strange mesonic sector. In the charmed baryonic sector, after departure from 1S1P-HRG, for the highest two temperatures, PDG-HRG prediction can very well describe $P^{C}_{B}$ in Fig. 2 [Left]. The fact that 1S1P-HRG lies above PDG-HRG for $P^{C}_{B}$ indicates that there are many low-lying and therefore thermodynamically dominant charmed baryons missing from the PDG record.
In the coexistence phase of charmed hadron and quark-like excitations, $m^{C}_{q}$ in Eq. (3) becomes temperature dependent and can be interpreted as the mass of a quasi-particle with quantum numbers of charm quark. Given $P^{C}_{q}(T)$ as a function of temperature, one can solve Eq. (3) at each temperature and obtain temperature dependence of $m^{C}_{q}$. In Fig. 3 [Right], at each temperature, the error on $m^{C}_{q}$ is the standard deviation of $m^{C}_{q}$ values obtained after solving Eq. (3) for $50$ fake Gaussian samples. Fig. 3 [Right] shows that at $T=162.4$ MeV, $m^{C}_{q}$ is around the mass of D-meson, and starts decreasing with temperature. Quasi-particle model in Eq. (4) considers a non-interacting gas of charmed hadron and quark-like excitations, but the temperature dependence of $m^{C}_{q}$ encodes the in-medium interactions. In the non-interacting quark gas limit, $m^{C}_{q}$ will become the pole mass of charm quark.